Lead is a poisonous metal that especially affects children because they retain a

larger fraction of lead than adults do. Lead levels of 0.250ppm in a child cause delayed

cognitive development. How many moles of lead present in 1.00g of child’s blood

would 0.250ppm represent.


1
Expert's answer
2021-10-29T02:20:31-0400

PPM is the measure of the concentration of contaminant present in a million parts of the solution.

Here, 0.250 PPM lead means 0.250 parts of lead present in one million parts of a child’s blood sample.

In terms of grams, 0.250 grams of lead in a blood sample.

Then, 1 gram of the sample contains,

grams of Lead in 1 g of blood sample "= \frac{0.250}{10^6} = 2.5 \times 10^{-7} \;g"

Obtain the moles of lead in 1 gram of sample by the use of the molar mass of lead as shown below:

"Moles\; of\; lead = \frac{mass}{molar \; mass} \\

= \frac{2.5 \times 10^{-7} \; g}{207.2 \; g/mol} \\

= 1.207 \times 10^{-9} \; moles"

Therefore, the moles of lead present in 1.00 g of child's blood is "1.207 \times 10^{-9} \;mol"